# quasimodular forms arising from Jacobi’s theta function and special symmetric polynomials

### v1

@@ Abstract.
Ramanujan derived a sequence of even weight $2n$ quasimodular forms $U_{2n}(q)$ from derivatives of Jacobi’s weight $3/2$ theta function. Using the generating function for this sequence, one can construct sequences of quasimodular forms of all nonnegative integer weights with minimal input: a weight 1 modular form and a power series $F(X)$. Using the weight 1 form $\theta(q)^{2}$ and $F(X)=e^{X/2}$, we obtain a sequence $\{Y_{n}(q)\}$ of weight $n$ quasimodular forms on $\Gamma_{0}(4)$ whose symmetric function avatars $\widetilde{Y}_{n}(\boldsymbol{x}^{k})$ are the symmetric polynomials $T_{n}(\boldsymbol{x}^{k})$ that arise naturally in the study of syzygies of numerical semigroups. With this information, we settle two conjectures about the $T_{n}(\boldsymbol{x}^{k}).$ Finally, we note that these polynomials are systematically given in terms of the Borel-Hirzebruch $\widehat{A}$-genus for spin manifolds, where one identifies power sum symmetric functions $p_{i}$ with Pontryagin classes.
@@ 1. Introduction and Statement of Results
Using the classical weight 3/2 Jacobi theta function [16, Thm. 1.60]
@@
Ramanujan (in his “lost notebook” [17, p. 369]) defined the sequence of $q$-series
@@
(1.3) $\theta(q)=\prod_{k\geq 1}(1-q^{2k})(1+q^{2k-1})^{2}=1+2q+2q^{4}+2q^{9}+2q^{16}+\dots$
@@
[…]
@@
${Y}_{n}(q)=\frac{1}{2^{n}(n+1)}\sum_{k=0}^{\lfloor\frac{n}{2}\rfloor}\binom{n+1}{2k+1}\theta(q)^{2n-4k}\operatorname{Tr}_{n}(\phi_{U};q).$
@@
(1.9) $\Psi(E_{\lambda})=\prod_{i=1}^{n}\Psi(E_{2i})^{m_{i}}=p_{2}(\boldsymbol{x})^{m_{1}}p_{4}(\boldsymbol{x})^{m_{2}}\cdots p_{2n}(\boldsymbol{x})^{m_{n}},$
@@
The second conjecture of [11] claimed striking identities. To state the conjecturecccWe note that we offer a more precise formulation of the conjecture [11, p. 62]., we require the numbers $A_{2j+1},$ known as the tangent/zig-zag numbers (counting up/down permutations), which have the exponential generating function
@@
These results illustrate the rich mathematics that arises from Ramanujan’s $U_{2n}(q)$ by letting $F(X)=e^{X/2}$ and $\alpha=1/2$ in (1.4). Hence, it is natural to ask the following open-ended questions.
@@ Acknowledgements
The authors thank Ajit Singh for his constructive comments on earlier versions of this paper. T.A. especially thanks Ira Gessel and Christophe Vignat for valuable discussions. The third author thanks the Thomas Jefferson Fund and the NSF (DMS-2002265 and DMS-2055118).
@@ 2.1. Symmetric functions arising from syzygies of numerical semigroups
[…]
where $\beta_{0}=1$, and
@@
and $C_{k,j}$ and $\beta_{k}$ stand for degrees of the $k$th syzygy and partial Betti’s numbers, respectively [20].
@@
where we let $\pi_{m}:=\prod_{i=1}^{m}d_{i}$. For $r\geq m,$ these sums were calculated and furnished in [10, (22)] as
@@
$\beta_{k}=\beta_{m-k-1},\quad\beta_{1}=m-1,\quad\beta_{m-1}=1,\quad C_{k,j}+C_{m-k-1,j}=C_{m-1,1},\quad\min\{d_{1},\dots,d_{m}\}\geq m.$
@@
(2.6) $H\left(S_{m}^{CI};z\right)=\frac{\prod_{j=1}^{m-1}\left(1-z^{e_{j}}\right)}{\prod_{i=1}^{m}\left(1-z^{d_{i}}\right)},\,\,e_{j}=c_{j}d_{j},\quad c_{j}\in{\mathbb{N}},\,\,c_{j}\geq 2,\,\,e_{j}\geq 2(m+1).$
The tuple $\boldsymbol{e}^{m-1}:=(e_{1},\ldots,e_{m-1})$ presents the $m-1$ degrees of the first syzygy for symmetric CI semigroup $S_{m}^{CI}$.
@@
[11, Lm. 1.3] We have the inequality $T_{n}(x_{1},\ldots,x_{m})\geq 0$ whenever $x_{1},\ldots,x_{m}\geq 0$.
@@
(2.9) $K_{t}(S_{m})=\sum_{r=0}^{t}{t\choose r}T_{t-r}(\sigma)G_{r}(S_{m})+\frac{2^{t+1}}{t+1}T_{t+1}(\delta),$
@@ 2.2. Symmetric functions arising from partitions
It is well-known (see, for example [9, Section 3.1]) that $W(s;\boldsymbol{d}^{k})$ can be described in terms of finitely many quasipolynomials $W_{q}(s;\boldsymbol{d}^{k})$, each containing a single $q$-periodic function with $W_{1}(s;\boldsymbol{d}^{k})$ being a polynomial, given in the form
@@
The $W_{q}(s;\boldsymbol{d}^{k})$’s are called the Sylvester waves. In particular, $W_{1}(s;\boldsymbol{d}^{k})$ is referred to as the first Sylvester wave and has fascinating properties. For example, $W_{1}(s;\boldsymbol{d}^{k})$ is the polynomial part of $W(s;\boldsymbol{d}^{k})$ having an explicit formula [9, (3.16) and (7.1)] given by
(2.11) $W_{1}(s;{\bf d}^{k})=\frac{1}{(k-1)!\pi_{k}}\sum_{j=0}^{k-1}{k-1\choose j}f_{j}({\bf d}^{k})s^{k-1-j},\qquad f_{j}({\bf d}^{k})=\left(\sigma_{1}+\sum_{i=1}^{k}{\mathcal{B}}d_{i}\right)^{j},$
where $\sigma_{1}:=\sum_{j=1}^{k}d_{j}$ and $\pi_{k}:=\prod_{i=1}^{k}d_{i}$. For our purpose in this paper, we will replace $\boldsymbol{d}^{k}=(d_{1},\dots,d_{k})$ by $\boldsymbol{x}^{k}=(x_{1},\dots,x_{k})$ and $\sigma_{j}$ by the power sum $p_{j}$ in the variables $(x_{1},\dots,x_{k})$. The Bernoulli symbol $\mathcal{B}$ is understood as $\mathcal{B}^{n}=B_{n}$ where
$\frac{z}{e^{z}-1}=\sum_{n\geq 0}B_{n}\cdot\frac{z^{n}}{n!}$
@@
(2.12) $f_{n}(\boldsymbol{x}^{k}):=\left(p_{1}+\sum_{i=1}^{k}\mathcal{B}x_{i}\right)^{n}=f_{n}(p_{1},\ldots,p_{n}),$
expanded in power sums (for example, see [9, (7.2)]), are:
@@
$f_{4}(\boldsymbol{x}^{k})$ $=\frac{15p_{1}^{4}-30p_{1}^{2}p_{2}+5p_{2}^{2}+2p_{4}}{240},\qquad f_{5}(\boldsymbol{x}^{k})=\frac{3p_{1}^{5}-10p_{1}^{2}p_{2}+5p_{1}p_{2}^{2}+2p_{1}p_{4}}{96},$
@@
The proof of Theorem 1.1 (1-2) follows easily from the Taylor expansion of $F(X)=e^{X/2}$ and direct algebraic manipulation. To obtain Theorem 1.2, one simply substitutes (1.8) (see also [1, Thm. 1.3 (1)]) into the formulas in Theorem 1.1 (2).
@@
If we let $\mathcal{B}_{n}:=B_{n}$ (Bernoulli numbers), $\mathcal{C}_{n}:=\frac{(-1)^{n}}{n+1}$ and $\mathcal{D}_{n}:=\frac{1}{n+1}$, then for positive integers $n$ in umbral formalism (treating $\mathcal{B},\mathcal{C}$ and $\mathcal{D}$ as random variables) the following are true:
@@
$e^{\mathcal{B}z}=\sum_{n\geq 0}\mathcal{B}^{n}\cdot\frac{z^{n}}{n!}=\sum_{n\geq 0}B_{n}\cdot\frac{z^{n}}{n!}=\frac{z}{e^{z}-1}.$
@@
$e^{(1+\mathcal{B})z}=\frac{ze^{z}}{e^{z}-1}=\frac{-z}{e^{-z}-1}=e^{(-\mathcal{B})z}.$
@@
$e^{\mathcal{D}z}=\sum_{n\geq 0}\mathcal{D}^{n}\frac{z^{n}}{n!}=\sum_{n\geq 0}\frac{1}{n+1}\frac{z^{n}}{n!}=\frac{e^{z}-1}{z}.$
@@
$e^{(\mathcal{B}+\mathcal{D})z}=\frac{z}{e^{z}-1}\cdot\frac{e^{z}-1}{z}=1=e^{0z},$
@@
$T_{n}(\boldsymbol{x}^{k})=\left(p_{1}+\sum_{i=1}^{k}\mathcal{C}\,x_{i}\right)^{n}.$
@@
Given a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$, let $m_{\lambda}=m_{\lambda_{1}}m_{\lambda_{2}}\cdots$ denote the associated monomial symmetric function [19, Chapter 7] by $m_{\lambda}$ defined by
@@
$\left(p_{1}+\sum_{i}\mathcal{C}x_{i}\right)^{n}=\left(\sum_{i}\mathcal{D}x_{i}\right)^{n}.$
@@
(3.1) $\left(\sum_{i=1}^{k}\mathcal{D}\,x_{i}\right)^{n}$ $=\sum_{\lambda\vdash n}\binom{n}{\lambda_{1},\lambda_{2},\dots}\mathcal{D}_{\lambda}m_{\lambda}.$
[…]
@@
$=\frac{(-1)^{k-1}(n+k)!\cdot\prod_{j=1}^{k}x_{j}}{n!}\left(\sum_{i=1}^{k}\mathcal{D}x_{i}\right)^{n},$
@@
The proof of Theorem 1.5 will also require the following fact about products of formal power series and the (unique) involution $\omega$ on symmetric functions, that is determined by the action $\omega(p_{n})=(-1)^{n-1}p_{n}$ on power sum functions (see, for example [19, Chapter 7]).
@@
$\left(p_{1}+\sum_{i}\mathcal{C}x_{i}\right)^{r}=\left(\sum_{i}\mathcal{D}x_{i}\right)^{r}\qquad\text{and}\qquad\left(p_{1}+\sum_{i}\mathcal{B}x_{i}\right)^{r}=\left(\sum_{i}(-\mathcal{B})x_{i}\right)^{r}.$
@@
$\left(\sum_{i}\mathcal{D}x_{i}\right)^{r}$ $=T_{r}(p_{1},p_{2},\dots)=\sum_{\lambda\vdash r}g^{\lambda}p_{\lambda}(\boldsymbol{x}),$
$\left(\sum_{i}\widetilde{\mathcal{B}}x_{i}\right)^{r}$ $=f_{r}(p_{1},p_{2},\dots)=\sum_{\lambda\vdash r}\widetilde{g}^{\lambda}p_{\lambda}(\boldsymbol{x}),$
for some coefficients $g^{\lambda}$ and $\widetilde{g}^{\lambda}$ where $\boldsymbol{x}=(x_{1},x_{2},\dots)$ and $\widetilde{\mathcal{B}}=-\mathcal{B}$.
@@
$f(z)=\frac{e^{z}-1}{z}\qquad{\text{\rm and}}\qquad g(z)=\frac{z}{1-e^{-z}},$
respectively, for $\mathcal{D}_{k}=\frac{1}{k+1}$ and $(-1)^{k}\mathcal{B}_{k}$. By extracting the degree $r$ terms (in $z$) and noting that $\omega(p_{k})=(-1)^{k-1}p_{k}$, we obtain the desired outcome that
@@
In this section, we make important use of Pólya’s cycle index theorem (PCIT) [19, Section 5]
@@
where we require writing a partition in the frequency notation $\lambda=(1^{m_{1}}\dots t^{m_{n}})\vdash n$. PCIT is deeply intertwined with other key results and methods, offering a consistent approach to a variety of counting problems that are similar and adjacent to the results in this paper. Specifically, it provides a method for deriving generating functions for structures related to generalized Bernoulli and Bell polynomials, Bernoulli-Barnes polynomials [6] (with interesting connection to Fourier Dedekind sums), and it forms a crucial part of the standard exponential generating functions due to Touchard. [19, Eq’n. (5.30)].
@@
We only prove the second assertion, as the first one follows a similarly. To this end, we recall the series expansion [13, 1.518.1] for the reciprocal of what we call here the sinhc function $\text{sinhc}(t)=\frac{\sinh(t)}{t}$ (reminiscent of the sinc function) given by
@@
$\exp(\mathcal{B}z)=\sum_{n\geq 0}\mathcal{B}^{n}\frac{z^{n}}{n!}=\sum_{n\geq 0}B_{n}\frac{z^{n}}{n!}=\frac{z}{e^{z}-1}$
@@
(3.4) $\exp((1+2\mathcal{B}_{i})x_{i}z)=\sum_{n\geq 0}((1+2\mathcal{B}_{i})x_{i})^{n}\frac{z^{n}}{n!}=\frac{2x_{i}z\cdot e^{x_{i}z}}{e^{2x_{i}z}-1}=\frac{1}{\text{sinhc}(x_{i}z)}.$
@@
Combining (3.3), (3.5) and Pólya’s cycle index theorem [19, Section 5], we arrive at
@@
In [11, (25)]), one of the authors proved the identity
@@
The use of umbral symbolism allows for the more compact reformulation
@@
$(f-f_{1})^{2n}$ $=\left(p_{1}+\sum_{i=1}^{k}\mathcal{B}_{i}x_{i}-\frac{1}{2}p_{1}\right)^{2n}=\frac{1}{4^{n}}\left(\sum_{i=1}^{k}(1+2\mathcal{B}_{i})x_{i}\right)^{2n}=\frac{\widetilde{u}_{2n}}{(2n+1)4^{n}},$
@@
In this subsection, we begin by recalling some calculations found in [11, p. 59, (13)], namely that
@@
The author [11] also derived a similar formulation for the polynomials $f_{n}(\boldsymbol{x}^{k})$. However, a compact formula is missing from [11, (25)-(26)] and we fill that gap. For integers $n\geq 1$, the modified version of [11, page 61, (25)] is

### v2

@@ Abstract.
Ramanujan derived a sequence of even weight $2n$ quasimodular forms $U_{2n}(q)$ from derivatives of Jacobi’s weight $3/2$ theta function. Using the generating function for this sequence, one can construct sequences of quasimodular forms of all nonnegative integer weights with minimal input: a weight 1 modular form and a power series $F(X)$. Using the weight 1 form $\theta(q)^{2}$ and $F(X)=\exp(X/2)$, we obtain a sequence $\{Y_{n}(q)\}$ of weight $n$ quasimodular forms on $\Gamma_{0}(4)$ whose symmetric function avatars $\widetilde{Y}_{n}(\boldsymbol{x}^{k})$ are the symmetric polynomials $T_{n}(\boldsymbol{x}^{k})$ that arise naturally in the study of syzygies of numerical semigroups. With this information, we settle two conjectures about the $T_{n}(\boldsymbol{x}^{k}).$ Finally, we note that these polynomials are systematically given in terms of the Borel-Hirzebruch $\widehat{A}-$genus for spin manifolds, where one identifies power sum symmetric functions $p_{i}$ with Pontryagin classes.
@@ 1. Introduction and Statement of Results
Using the classical weight 3/2 Jacobi theta function [19, Thm. 1.60]
@@
Ramanujan (in his “lost notebook” [20, p. 369]) defined the sequence of $q$-series
@@
(1.3) $\theta(q)=\prod_{k\geq 1}(1-q^{2k})(1+q^{2k-1})^{2}=1+2q+2q^{4}+2q^{9}+2q^{16}+\cdots.$
@@
$\phi_{U}(\lambda):=(2n+1)!\,4^{n}\cdot\prod_{j=1}^{n}\frac{1}{m_{j}!}\left(\frac{B_{2j}}{(2j)(2j)!}\right)^{m_{j}}.$
@@
${Y}_{n}(q)=\frac{1}{2^{n}(n+1)}\sum_{k=0}^{\lfloor\frac{n}{2}\rfloor}\binom{n+1}{2k+1}\theta(q)^{2n-4k}\operatorname{Tr}_{k}(\phi_{U};q).$
Remark.
We could have declared $\theta(q)^{2}$ as an Eisenstein series allowing us to reformulate the sums in Theorem 1.2 as partition Eisenstein traces over $\{\theta(q)^{2},E_{2},E_{4},E_{6},\dots\}$, along the lines of recent work by Bringmann, Pandey, and van Ittersum [9]. However, this would have required unnecessary new notation.
@@
(1.9) $\Psi(E_{\lambda})=\prod_{i=1}^{n}\Psi(E_{2i})^{m_{i}}=p_{\lambda}(\boldsymbol{x})\qquad\text{where}\qquad p_{\lambda}(\boldsymbol{x}):=p_{2}(\boldsymbol{x})^{m_{1}}p_{4}(\boldsymbol{x})^{m_{2}}\cdots p_{2n}(\boldsymbol{x})^{m_{n}},$
@@
The second conjecture of [12] claimed striking identities. To state the conjecturecccWe note that we offer a more precise formulation of the conjecture [12, p. 62]., we require the numbers $A_{2j+1},$ known as the tangent/zig-zag numbers (counting up/down permutations), which have the exponential generating function
@@
These results illustrate the rich mathematics that arises from Ramanujan’s $U_{2n}(q)$ by letting $F(X)=\exp(X/2)$ and $\alpha=1/2$ in (1.4). Hence, it is natural to ask the following open-ended questions.
@@ Acknowledgements
The authors thank Toshiki Matsusaka, Badri Pandey, Ajit Singh for their constructive comments on earlier versions of this paper. T.A. especially thanks Ira Gessel and Christophe Vignat for valuable discussions. The third author thanks the Thomas Jefferson Fund and the NSF (DMS-2002265 and DMS-2055118).
@@ 2.1. Symmetric functions arising from syzygies of numerical semigroups
(2.2) $H\left(S_{m};z\right)=\frac{Q\left(S_{m};z\right)}{\prod_{i=1}^{m}\left(1-z^{d_{i}}\right)},$
where
@@
while $C_{k,j}\in{\mathbb{N}},\;\;1\leq k\leq m-1,\;\;1\leq j\leq\beta_{k}$, and $C_{k,j}$ and $\beta_{k}$ stand for degrees of the $k$th syzygy and partial Betti’s numbers, respectively [23]. Here $\beta_{0}=1$.
@@
where we let $\pi_{m}:=\prod_{i=1}^{m}d_{i}$. For $r\geq m,$ these sums were calculated and furnished in [11, (22)] as
@@
$\beta_{k}=\beta_{m-k-1},\quad\beta_{1}=m-1,\quad\beta_{m-1}=1,\quad C_{k,j}+C_{m-k-1,j}=C_{m-1,1},\quad\min\{d_{1},\dots,d_{m}\}\geq m+1.$
@@
(2.6) $H\left(S_{m}^{CI};z\right)=\frac{\prod_{j=1}^{m-1}\left(1-z^{e_{j}}\right)}{\prod_{i=1}^{m}\left(1-z^{d_{i}}\right)}\qquad\text{and}\qquad e_{j}\geq 2(m+1).$
The tuple $\boldsymbol{e}^{m-1}:=(e_{1},\ldots,e_{m-1})$ presents the m1m-1 degrees of the first syzygy for symmetric CI semigroup $S_{m}^{CI}$.
@@
[12, Lm. 1.3] We have the inequality $T_{n}(x_{1},\ldots,x_{m})\geq 0$ whenever $x_{1},\ldots,x_{m}\geq 0$.
@@
(2.9) $K_{t}(S_{m})=\sum_{r=0}^{t}\binom{t}{r}T_{t-r}(\sigma)G_{r}(S_{m})+\frac{2^{t+1}}{t+1}T_{t+1}(\delta),$
@@ 2.2. Symmetric functions arising from partitions
It is well-known (see, for example [10, Section 3.1]) that $W(s;\boldsymbol{d}^{k})$ can be described in terms of finitely many quasipolynomials $W_{q}(s;\boldsymbol{d}^{k})$, each containing a single $q$-periodic function with $W_{1}(s;\boldsymbol{d}^{k})$ being a polynomial, given in the form
@@
The $W_{q}(s;\boldsymbol{d}^{k})$’s are called the Sylvester waves. In particular, $W_{1}(s;\boldsymbol{d}^{k})$ is referred to as the first Sylvester wave and has fascinating properties. For example, $W_{1}(s;\boldsymbol{d}^{k})$ is the polynomial part of $W(s;\boldsymbol{d}^{k})$ having an explicit formula [10, (3.16) and (7.1)] given by
(2.11) $W_{1}(s;{\bf d}^{k})=\frac{1}{(k-1)!\pi_{k}}\sum_{j=0}^{k-1}\binom{k-1}{j}f_{j}({\bf d}^{k})s^{k-1-j},\qquad f_{j}({\bf d}^{k})=\left(\sigma_{1}+\sum_{i=1}^{k}\mathcal{B}_{i}d_{i}\right)^{j},$
where $\sigma_{1}:=\sum_{j=1}^{k}d_{j}$ and $\pi_{k}:=\prod_{i=1}^{k}d_{i}$. For our purpose in this paper, we will replace $\boldsymbol{d}^{k}=(d_{1},\dots,d_{k})$ by $\boldsymbol{x}^{k}=(x_{1},\dots,x_{k})$ and $\sigma_{j}$ by the power sum $p_{j}$ in the variables $(x_{1},\dots,x_{k})$. By treating the $\mathcal{B}_{i}$’s as independent random variables, the Bernoulli symbolism $(\mathcal{B}_{i}x_{i})^{n}$ is understood as $\mathcal{B}_{i}^{n}x_{i}=B_{i}^{n}x_{i}^{n}$ where
$\frac{z}{\exp(z)-1}=\sum_{n\geq 0}B_{n}\cdot\frac{z^{n}}{n!}$
@@
(2.12) $f_{n}(\boldsymbol{x}^{k}):=\left(p_{1}+\sum_{i=1}^{k}\mathcal{B}_{i}x_{i}\right)^{n}=f_{n}(p_{1},\ldots,p_{n}),$
expanded in power sums (for example, see [10, (7.2)]), are:
@@
$f_{4}(\boldsymbol{x}^{k})$ $=\frac{15p_{1}^{4}-30p_{1}^{2}p_{2}+5p_{2}^{2}+2p_{4}}{240},\qquad f_{5}(\boldsymbol{x}^{k})=\frac{3p_{1}^{5}-10p_{1}^{3}p_{2}+5p_{1}p_{2}^{2}+2p_{1}p_{4}}{96},$
@@
Remark.
If we set all of the variables to $x$, then we obtain the convolution Stirling polynomials studied by Knuth [17]
$f_{1}(x)=\frac{x}{2},\ \ f_{2}(x,x)=\frac{3x^{2}-x}{12},\ \ f_{3}(x,x,x)=\frac{x^{3}-x^{2}}{8},\ \ f_{4}(x,\dots,x)=\frac{15x^{4}-30x^{3}+5x^{2}+2x}{240}\dots,$
which has generating function
$\left(\frac{z\exp(z)}{\exp(z)-1}\right)^{x}=\sum_{n=0}^{\infty}f_{n}(x,\dots,x)\cdot\frac{z^{n}}{n!}.$
@@
The proof of Theorem 1.1 (1-2) follows easily from the Taylor expansion of $F(X)=\exp(X/2)$ and direct algebraic manipulation. To obtain Theorem 1.2, one simply substitutes (1.8) (see also [1, Thm. 1.3 (1)]) into the formulas in Theorem 1.1 (2).
@@
If we let $\mathcal{B}^{n}=B_{n}$ (Bernoulli numbers), $\mathcal{C}_{n}:=\frac{(-1)^{n}}{n+1}$ and $\mathcal{D}_{n}:=\frac{1}{n+1}$, then for positive integers $n$ in umbral formalism (treating $\mathcal{B},\mathcal{C}$ and $\mathcal{D}$ as random variables) the following are true:
@@
$\exp(\mathcal{B}z)=\sum_{n\geq 0}\mathcal{B}^{n}\cdot\frac{z^{n}}{n!}=\sum_{n\geq 0}B_{n}\cdot\frac{z^{n}}{n!}=\frac{z}{\exp(z)-1}.$
@@
$\exp((1+\mathcal{B})z)=\frac{z\exp(z)}{\exp(z)-1}=\frac{-z}{\exp(-z)-1}=\exp((-\mathcal{B})z).$
@@
$\exp(\mathcal{D}z)=\sum_{n\geq 0}\mathcal{D}^{n}\frac{z^{n}}{n!}=\sum_{n\geq 0}\frac{1}{n+1}\frac{z^{n}}{n!}=\frac{\exp(z)-1}{z}.$
@@
$\exp((\mathcal{B}+\mathcal{D})z)=\frac{z}{\exp(z)-1}\cdot\frac{\exp(z)-1}{z}=1=\exp(0z),$
@@
$T_{n}(\boldsymbol{x}^{k})=\left(p_{1}+\sum_{i=1}^{k}\mathcal{C}_{i}x_{i}\right)^{n}.$
@@
Given a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$, let $m_{\lambda}$ denote the associated monomial symmetric function [22, Chapter 7] by $m_{\lambda}$ defined by
@@
$\left(p_{1}+\sum_{i}\mathcal{C}_{i}x_{i}\right)^{n}=\left(\sum_{i}\mathcal{D}_{i}x_{i}\right)^{n}.$
@@
(3.1) $\left(\sum_{i=1}^{k}\mathcal{D}_{i}x_{i}\right)^{n}$ $=\sum_{\lambda\vdash n}\binom{n}{\lambda_{1},\lambda_{2},\dots}\mathcal{D}_{\lambda}m_{\lambda}.$
We apply (3.1) with $\mathcal{D}$ replaced by $\boldsymbol{1}=(1,1,\dots)$ and $n$ replaced by $n^{\prime}:=n+k$, while we rank partitions according to their lengths, to reformulate (2.7) to obtain
@@
$=\frac{(-1)^{k-1}(n+k)!\cdot\prod_{j=1}^{k}x_{j}}{n!}\left(\sum_{i=1}^{k}\mathcal{D}_{i}x_{i}\right)^{n},$
@@
The proof of Theorem 1.5 will also require the following fact about products of formal power series and the (unique) involution $\omega$ on symmetric functions, that is determined by the action $\omega(p_{n})=(-1)^{n-1}p_{n}$ on power sum functions (see, for example [22, Chapter 7]).
@@
$\left(p_{1}+\sum_{i}\mathcal{C}_{i}x_{i}\right)^{r}=\left(\sum_{i}\mathcal{D}_{i}x_{i}\right)^{r}\qquad\text{and}\qquad\left(p_{1}+\sum_{i}\mathcal{B}_{i}x_{i}\right)^{r}=\left(\sum_{i}(-\mathcal{B}_{i})x_{i}\right)^{r}.$
@@
$\left(\sum_{i}\mathcal{D}_{i}x_{i}\right)^{r}$ $=T_{r}(p_{1},p_{2},\dots)=\sum_{\lambda\vdash r}g^{\lambda}p_{\lambda}(\boldsymbol{x}),$
$\left(\sum_{i}\widetilde{\mathcal{B}}_{i}x_{i}\right)^{r}$ $=f_{r}(p_{1},p_{2},\dots)=\sum_{\lambda\vdash r}\widetilde{g}^{\lambda}p_{\lambda}(\boldsymbol{x}),$
for some coefficients $g^{\lambda}$ and $\widetilde{g}^{\lambda}$ where $\boldsymbol{x}=(x_{1},x_{2},\dots)$ and $\widetilde{\mathcal{B}}_{i}=-\mathcal{B}_{i}$.
@@
$f(z)=\frac{\exp(z)-1}{z}\qquad{\text{\rm and}}\qquad g(z)=\frac{z}{1-\exp(-z)},$
respectively, for $\mathcal{D}_{k}=\frac{1}{k+1}$ and $(-1)^{k}B_{k}$. By extracting the degree $r$ terms (in $z$) and noting that $\omega(p_{k})=(-1)^{k-1}p_{k}$, we obtain the desired outcome that
@@
In this section, we make important use of Pólya’s cycle index theorem (PCIT) [22, Section 5]
@@
where we require writing a partition in the frequency notation $\lambda=(1^{m_{1}}\dots t^{m_{n}})\vdash n$. PCIT is deeply intertwined with other key results and methods, offering a consistent approach to a variety of counting problems that are similar and adjacent to the results in this paper. Specifically, it provides a method for deriving generating functions for structures related to generalized Bernoulli and Bell polynomials [18], Bernoulli-Barnes polynomials [6] (with interesting connection to Fourier Dedekind sums), and it forms a crucial part of the standard exponential generating functions due to Touchard [22, Eq’n. (5.30)].
@@
We only prove the second assertion, as the first one follows a similarly. To this end, we recall the series expansion [14, 1.518.1] for the reciprocal of what we call here the sinhc function $\text{sinhc}(t)=\frac{\sinh(t)}{t}$ (reminiscent of the sinc function) given by
@@
$\exp(\mathcal{B}z)=\sum_{n\geq 0}\mathcal{B}^{n}\frac{z^{n}}{n!}=\sum_{n\geq 0}B_{n}\frac{z^{n}}{n!}=\frac{z}{\exp(z)-1}$
@@
(3.4) $\exp((1+2\mathcal{B}_{i})x_{i}z)=\sum_{n\geq 0}((1+2\mathcal{B}_{i})x_{i})^{n}\frac{z^{n}}{n!}=\frac{2x_{i}z\cdot\exp(x_{i}z)}{\exp(2x_{i}z)-1}=\frac{1}{\text{sinhc}(x_{i}z)}.$
@@
Combining (3.3), (3.5) and Pólya’s cycle index theorem [22, Section 5], we arrive at
@@
In [12, (25)]), one of the authors proved the identity
@@
Letting $\mathfrak{f}:=p_{1}+\sum_{i=1}^{k}\mathcal{B}_{i}x_{i}$ The use of umbral symbolism allows for the more compact reformulation
@@
$(\mathfrak{f}-f_{1})^{2n}$ $=\left(p_{1}+\sum_{i=1}^{k}\mathcal{B}_{i}x_{i}-\frac{1}{2}p_{1}\right)^{2n}=\frac{1}{4^{n}}\left(\sum_{i=1}^{k}(1+2\mathcal{B}_{i})x_{i}\right)^{2n}=\frac{\widetilde{u}_{2n}}{(2n+1)4^{n}},$
@@
In this subsection, we begin by recalling some calculations found in [12, p. 59, (13)], namely that
@@
The author [12] also derived a similar formulation for the polynomials $f_{n}(\boldsymbol{x}^{k})$. However, a compact formula is missing from [12, (25)-(26)] and we fill that gap. For integers $n\geq 1$, the modified version of [12, page 61, (25)] is